Solve each equation by completing the square. If necessary round to the nearest hundredth.
a^2+ 8a=513

Answer by Cromlix(4381)

Hi there,
a^2 + 8a = 513
(a + 4)^2 = (-4)^2 + 513
(a + 4)^2 = 529
a+4 = +/- √529
a = -4 +/- √529
a = -4 - 23 or a = -4 + 23
a = - 27 or a = 19
Hope this helps :-)